you observe the last dividend paid of $4 per share. The growth rate is projected to be a constant 5% per year. Your required rate of return is 8%.
a. what price are you willing to pay for that stock today?
b. what is expected stock price in 1 year?
c. find dividend yield
d. find capital gains
Answer #1

Pirce today = D1 = (ke- g) $140.00 (4*1.05)/(0.08-0.05) Price after 1 year = $147.00 (140*1.05] non Dividend yield = dividend